A motorcycle that travels 80kph starts 45 minutes after a pick-up truck and overtakes it in 15 minutes. Find the speed of the truck.
Answer by mananth(16946)

motorcycle = 80kph
starts 45 minutes after a pick-up truck 3/4 hours
overtakes it in 15 minutes. = 1/4 hour
let speed of truck = x kmph
time traveled by truck = 1 hour
distance traveled by truck = x miles in 1 hour
..
distance traveled by motorcycle = 80*3/4 = 60 miles
both travel same distance when motorcycle catches up.
so truck traveled 60 miles
..
time 1 hour distance 60 km
speed of truck = 60km/h
...
m.ananth@hotmail.ca